OpenStudy (anonymous):

If a^2+ab+b^2=0 ,Prove that Log(a+b)=1/2 (loga+logb)

OpenStudy (anonymous):

I'am just checking da answer .....

OpenStudy (anonymous):

OH sure, why don't you add your work then?

OpenStudy (anonymous):

\[(a+b)^2=a^2+2ab+b^2=ab\] is a start

OpenStudy (anonymous):

should work neatly from there

OpenStudy (anonymous):

\[ a+b = \sqrt{ab} \implies \log (a+b) =\frac 12 \log (ab) = \frac 12 (\log a+\log b) \]

OpenStudy (anonymous):

here is it : ____________ a^2+ab+b^2=0 , (add ab and subtract ab) a^2+2ab+b^2=ab ,(a+b)^2=ab ..Then take log the da both sides log(a+b)^2=log(ab) ,2log(a+b)=log(ab) (divide by 2) log(a+b)=1/2(loga+logb) _____________________________________________________________________________ Check this out @FoolForMath @satellite73

OpenStudy (anonymous):

am i doing right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good ty